# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

### v2

@@ Abstract
Oversmoothing is a fundamental challenge in graph neural networks (GNNs): as the number of layers increases, node embeddings become increasingly similar, and model performance drops sharply. Traditionally, oversmoothing has been quantified using metrics that measure the similarity of neighbouring node features, such as the Dirichlet energy. While these metrics are related to oversmoothing, we argue they have critical limitations and fail to reliably capture oversmoothing in realistic scenarios. For instance, they provide meaningful insights only for very deep networks and under somewhat strict conditions on the norm of network weights and feature representations. As an alternative, we propose measuring oversmoothing by examining the numerical or effective rank of the feature representations. We provide theoretical support for this approach, demonstrating that the numerical rank of feature representations converges to one for a broad family of nonlinear activation functions under the assumption of nonnegative trained weights. To the best of our knowledge, this is the first result that proves the occurrence of oversmoothing without assumptions on the boundedness of the weight matrices. Along with the theoretical findings, we provide extensive numerical evaluation across diverse graph architectures. Our results show that rank-based metrics consistently capture oversmoothing, whereas energy-based metrics often fail. Notably, we reveal that a significant drop in the rank aligns closely with performance degradation, even in scenarios where energy metrics remain unchanged.
Machine Learning, ICML
@@ 1 Introduction
Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning (Peng et al., 2023; Tian et al., 2022), personalised recommendation systems (Damianou et al., 2024; Peng et al., 2022), social network analysis (Fan et al., 2019), and 3D mesh classification (Shi & Rajkumar, 2020). Central to most GNN architectures is the message-passing paradigm, where node features are iteratively aggregated from their neighbours and transformed using learned functions, such as multi-layer perceptrons or graph-attention mechanisms.
However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a soft limit on the depth of GNNs. This is often linked to the observation of increasing similarity between learned features as GNNs deepen and is named oversmoothing (Li et al., 2018; Nt & Maehara, 2019).
Over the years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants (Oono & Suzuki, 2019; Cai & Wang, 2020; Bodnar et al., 2022; Nguyen et al., 2022; Di Giovanni et al., 2023; Wu et al., 2023; Roth & Liebig, 2023). At a high level, most of these metrics directly measure the norm of the absolute deviation from the dominant eigenspace of the message-passing matrix. In linear GNNs without bias terms, this eigenspace is often known and easily computable via e.g. the power method. However, when nonlinear activation functions or biases are used, the dominant eigenspace may change, causing these oversmoothing metrics to fail and give false negative signals about the oversmoothing state of the learned features.
Therefore, these metrics are often considered as providing sufficient but not necessary evidence for oversmoothing (Rusch et al., 2023a). Despite this, there is a considerable body of literature using these somewhat unreliable metrics as part of their evidence for non-occurrence of oversmoothing in GNNs (Zhou et al., 2021; Chen et al., 2022; Rusch et al., 2022; Wang et al., 2022; Maskey et al., 2023; Nguyen et al., 2023; Rusch et al., 2023b; Epping et al., 2024; Scholkemper et al., 2024; Wang & Cho, 2024).
As we show in Section 6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed. Most empirical studies in the literature that observe the decay of the Dirichlet-energy-like metrics are conducted over the layers of the same very deep untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable. GNNs (Rusch et al., 2022; Wang et al., 2022; Rusch et al., 2023b; Wu et al., 2023), where the decay of the metrics is likely driven by the small weight initializations reducing all features to zero. Instead, we observe that when GNNs of different depths are separately trained, these metrics do not correlate well with their performance degradation.
@@
  • We review popular oversmoothing metrics in the current literature and provide a novel perspective on their theoretical analysis based on nonlinear activation eigenvectors.
  • Based on this analysis and observations on the linear case, we propose the rank as a better metric for quantifying oversmoothing, and thereby re-defining oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
  • We prove the convergence of the numerical rank towards one for linear GNNs and nonlinear GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function under the assumption of non-negative weights. To our knowledge, this is the first theoretical result proving that oversmoothing may occur independently of the weights’ magnitude with nonlinear activation functions.
@@ 2.1 Graph Convolutional Network
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $\mathcal{V}$ denoting its set of vertices and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ its set of edges. Let $\widetilde{A}\in\mathbb{R}^{N\times N}$ be the unweighted adjacency matrix with $N=|\mathcal{V}|$ being the total number of nodes, $|\mathcal{E}|$ being the total number of edges of $\mathcal{G}$ and $A$ the corresponding symmetric adjacency matrix normalized by the node degrees:
$A=\widetilde{D}^{-1/2}\widetilde{A}\widetilde{D}^{-1/2},$ (1)
where $\widetilde{D}=D+I$, $D$ is the diagonal degree matrix of the graph $\mathcal{G}$, and $I$ is the identity matrix. The rows of the feature matrix $X\in\mathbb{R}^{N\times d}$ are the concatenation of the $d$-dimensional feature vectors of all nodes in the graph. At each layer $l$, the GCN updates the node features as follows
$X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ (2)
where $\sigma$ is a nonlinear activation function, applied componentwise, and $W^{(l)}$ is a trainable weight matrix.
@@ 2.2 Graph Attention Network
While GCNs use a fixed normalized adjacency matrix to perform graph convolutions at each layer, Graph Attention Networks (GATs) (Veličković et al., 2017; Brody et al., 2021) perform graph convolution through a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism as follows
$A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i}+p_{2}^{(l)\top}W^{(l)\top}X_{j}))$ (3)
where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i},X_{j}$ denote the feature of the $i$th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization
$\text{softmax}_{j}(A_{ij})=\frac{\exp(A_{ij})}{\sum_{j^{\prime}}\exp(A_{ij^{\prime}})}.$ (4)
The corresponding feature update is
$X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$ (5)
@@ 3 Oversmoothing
Oversmoothing can be broadly understood as an increase in similarity between node features as inputs are propagated through an increasing number of message-passing layers, leading to a noticeable decline in GNN performance. However, the precise definition of this phenomenon varies across different sources. Some works define oversmoothing more rigorously as the alignment of all feature vectors with each other. This definition is motivated by the behaviour of a linear GCN:
$X^{(l+1)}=A\cdots AX^{(0)}W^{(0)}\dots W^{(l)}.$ (6)
Indeed if $\widetilde{A}$ is the adjacency matrix of a fully connected graph $\mathcal{G}$, $A$ will have a spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will thus converge towards the eigenspace spanned by the dominant eigenvector. Precisely, we have
$A^{l}\to uv^{\top}\quad\text{as}\quad l\rightarrow\infty,$ (7)
where $Au=u$, and $A^{\top}v=v$, see e.g. (Tudisco et al., 2015).
As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ converges in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$. Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{T}X^{(l)}\|$, converges to zero.
@@ 3.1 Existing Oversmoothing Metrics
Motivated by the discussion about the linear case, oversmoothing is thus quantified and analysed in terms of the convergence of some node similarity metrics towards zero. In particular, in most cases, it is measured exactly by the alignment of the features with the dominant eigenvector of the matrix $A$. The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features (Cai & Wang, 2020; Rusch et al., 2023a)
$E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i}}{u_{i}}-\frac{X_{j}}{u_{j}}\right\|^{2}_{2},$ (8)
where $u_{i}$ is the $i$-th entry of the dominant eigenvector of the message-passing matrix in (1). It thus immediately follows from our discussion on the linear setting that $E_{\mathrm{Dir}}(X^{(l)})$ converges to zero as $l\to\infty$ for a linear GCN with converging weights product $W^{(0)}\cdots W^{(l)}$. This intuition suggests that a similar behaviour may occur for “smooth-enough” nonlinearities.
In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and (Cai & Wang, 2020) have proved that, using LeakyReLU activation functions $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
Similarly, in the case of GATs, the matrices $A_{i}$ are all row stochastic, meaning that $u_{i}=1$ for all $i$. In this case, in (Wu et al., 2023), it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero
@@ 3.2 A unifying perspective based on the eigenvectors of nonlinear activations
We say that a vector $u\in\mathbb{R}^{N}\setminus\{0\}$ is an eigenvector of the (nonlinear) activation function $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{n}$ if for any $t\in\mathbb{R}\setminus\{0\}$, there exists $\mu_{t}$ such that $\sigma(tu)=\mu_{t}u$.
@@
This assumption recurs throughout our theoretical analysis, aligning with existing results in the literature. For example, (a) the proof by Cai & Wang (2020) applies to GCNs with LeakyReLU, where the dominant eigenvector of $A$ is nonnegative by the Perron-Frobenius theorem and, therefore, an eigenvector of LeakyReLU; and (b) the proof by Wu et al. (2023) holds for stochastic message-passing matrices $A^{(l)}$, which inherently share a common dominant eigenvector with constant entries, that is an eigenvector for any pointwise nonlinear activation function.
Lemma 3.2.
Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ be the $l$-th layer of a GNN. Assume that $A^{(l)}$ has a simple dominant eigenvalue with corresponding eigenvector $u$ that is also an eigenvector of $\sigma$. If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, then
$\|(I-\mathcal{P})X^{(l+1)}\|_{F}\leq\|(I-\mathcal{P})A^{(l)}X^{(l)}W^{(l)}\|_{F},$
where $\mathcal{P}=uu^{T}/\|u\|^{2}$ is the projection matrix on the linear space spanned by $u$.
The following theorem follows as a consequence.
@@
With the same notation of Lemma 3.2, if
$\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0,$
then $E_{\mathrm{Proj}}(X^{(l)})\to 0$ as $l\to\infty$.
Proof.
The proof follows observing that in the decomposition $(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}\mathcal{P}+(I-\mathcal{P})A^{(l)}(I-\mathcal{P})$ the matrix $(I-\mathcal{P})A^{(l)}\mathcal{P}$ is zero because $A^{(l)}u=\lambda_{1}^{l}u$ for any $l$. Thus
$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
and, from Lemma 3.2 and the inequality $\|AB\|_{F}\leq\|A\|_{2}\|B\|_{F}$, we have $\|(I-\mathcal{P})X^{(L)}\|_{F}\leq\Big{(}\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}\Big{)}\|X^{(0)}\|_{F}$. ∎
Note, in particular, that in the case of GCNs, the matrix $A^{(l)}=A$ is symmetric, and thus $\|I-\mathcal{P}A^{(l)}\|_{2}=\lambda_{2}$. Therefore, when $\sigma=\text{LeakyReLU}$, we obtain the result by Cai & Wang (2020) as convergence to zero is guaranteed if $\|W^{(l)}\|_{2}\leq\lambda_{2}$. Recall that, as discussed above, the choice $\sigma=\text{LeakyReLU}$ satisfies our eigenvector assumption since $u\geq 0$ by the Perron-Frobenius theorem, and thus LeakyReLU$(tu)=\alpha tu$ with $\alpha$ depending only on the sign of $t$. Similarly, in the case of GATs, the matrices $A^{(l)}$ are stochastic for all $l$, implying that $u=\mathbbm{1}$ is the constant vector with $(u)_{i}=1$ for all $i$. If $\sigma=\otimes\psi$ is a nonlinear activation function acting entrywise through $\psi$, then $\sigma(t\mathbbm{1})=\psi(t)\mathbbm{1}$. Therefore, Theorem 3.3 implies that if the weights are sufficiently small, the features align independently of the activation function used. This is consistent with the results in (Wu et al., 2023). However, we note that the bounds on the weights required by Theorem 3.3 and those in (Wu et al., 2023) on the weights $W^{(l)}$ are not identical, and it is unclear which of the two is more significant. Nonetheless, in both cases, having bounded weights along with any $1$-Lipschitz pointwise activation function is a sufficient condition for observing oversmoothing via $E_{\mathrm{Proj}}$ in GATs. In addition to offering a different and unifying theoretical perspective on the results in (Cai & Wang, 2020; Wu et al., 2023), we highlight the simplicity of our eigenvector-based proof, which we hope provides added clarity on this phenomenon.

### v3

@@ Abstract
Oversmoothing is a fundamental challenge in graph neural networks (GNNs): as the number of layers increases, node embeddings become increasingly similar, and model performance drops sharply. Traditionally, oversmoothing has been quantified using metrics that measure the similarity of neighbouring node features, such as the Dirichlet energy. While these metrics are related to oversmoothing, we argue they have critical limitations and fail to reliably capture oversmoothing in realistic scenarios. For instance, they provide meaningful insights only for very deep networks and under somewhat strict conditions on the norm of network weights and feature representations. As an alternative, we propose measuring oversmoothing by examining the numerical or effective rank of the feature representations. We provide theoretical support for this approach, demonstrating that the numerical rank of feature representations converges to one for a broad family of nonlinear activation functions under the assumption of nonnegative trained weights. To the best of our knowledge, this is the first result that proves the occurrence of oversmoothing in the nonlinear setting without assumptions on the boundedness of the weight matrices. Along with the theoretical findings, we provide extensive numerical evaluation across diverse graph architectures. Our results show that rank-based metrics consistently capture oversmoothing, whereas energy-based metrics often fail. Notably, we reveal that a significant drop in the rank aligns closely with performance degradation, even in scenarios where energy metrics remain unchanged.
@@ 1 Introduction
Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning [25, 39], personalised recommendation systems [7, 26], social network analysis [10], and 3D mesh classification [37]. Central to most GNN architectures is the message-passing paradigm, where node features are iteratively aggregated from their neighbours and transformed using learned functions, such as multi-layer perceptrons or graph-attention mechanisms.
However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a limit on their depth. This issue, often linked to the increasingly similar learned features as GNNs deepen, is known as oversmoothing [19, 23].
Over the years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants [24, 4, 2, 21, 8, 45, 30]. At a high level, most of these metrics directly measure the norm of the absolute deviation from the dominant eigenspace of the message-passing matrix. In linear GNNs without bias terms, this eigenspace is often known and easily computable via e.g. the power method. However, when nonlinear activation functions or biases are used, the dominant eigenspace may change, causing these oversmoothing metrics to fail and give false negative signals about the oversmoothing state of the learned features.
Therefore, these metrics are often considered as sufficient but not necessary evidences for oversmoothing [33]. Despite this, there is a considerable body of literature using these unreliable metrics as their evidences for non-occurrence of oversmoothing in GNNs [47, 6, 35, 44, 20, 22, 34, 9, 36, 43, 29, 42].
As we show in Section 6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed. Most empirical studies in the literature that observe the decay of the Dirichlet-energy-like metrics are conducted over the layers of the same very deep untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable. GNNs [44, 35, 34, 45, 29, 42], where the decay of the metrics is likely driven by the small weight initializations reducing all features to zero. Instead, we observe that when GNNs of different depths are separately trained, these metrics do not correlate well with their performance degradation.
@@
  • We review popular oversmoothing metrics in the current literature and simplify their theoretical analysis from a novel perspective of nonlinear activation eigenvectors.
  • We propose the rank as a better metric for quantifying oversmoothing, and thereby re-defining oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
  • We prove the convergence of the numerical rank towards one for linear GNNs and nonlinear GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function under the assumption of non-negative weights. To the best of our knowledge, this is the first theoretical result proving that oversmoothing may occur independently of the weights’ magnitude with nonlinear activation functions.
@@ 2.1 Graph Convolutional Network
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $\mathcal{V}$ denoting its set of vertices and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ its set of edges. Let $\widetilde{A}\in\mathbb{R}^{N\times N}$ be the unweighted adjacency matrix with $N=|\mathcal{V}|$ being the total number of nodes, $|\mathcal{E}|$ being the total number of edges of $\mathcal{G}$ and $A$ the corresponding symmetric adjacency matrix normalized by the node degrees: $A=\widetilde{D}^{-1/2}\widetilde{A}\widetilde{D}^{-1/2}$, where $\widetilde{D}=D+I$, $D$ is the diagonal degree matrix of the graph $\mathcal{G}$, and $I$ is the identity matrix. The rows of the feature matrix $X\in\mathbb{R}^{N\times d}$ are the concatenation of the $d$-dimensional feature vectors of all nodes in the graph. At each layer $l$, the node feature update of Graph Convolutional Network (GCN) [17] follows $X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ where $\sigma$ is a nonlinear activation function, applied component-wise, and $W^{(l)}$ is a trainable weight matrix.
@@ 2.2 Graph Attention Network
Graph Attention Networks (GATs) [41, 3] perform graph convolution via a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism $A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i,:}+p_{2}^{(l)\top}W^{(l)\top}X_{j,:}))$ where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i,:},X_{j,:}$ denote the feature of the $i$-th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization $\text{softmax}_{j}(A_{ij})=\exp(A_{ij})/\sum_{j^{\prime}}\exp(A_{ij^{\prime}})$. The corresponding feature update is $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$
@@ 3 Oversmoothing
Oversmoothing can be broadly understood as an increase in similarity between node features as inputs are propagated through an increasing number of message-passing layers, leading to a noticeable decline in GNN performance. However, the precise definition of this phenomenon varies across different sources. Some works define oversmoothing more rigorously as the alignment of all feature vectors with each other. This definition is motivated by the behaviour of a linear GCN: $X^{(l+1)}=A\cdots AX^{(0)}W^{(0)}\dots W^{(l)}$. Indeed, if $\widetilde{A}$ is the adjacency matrix of a fully connected graph, $A$ will have spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will converge toward the eigenspace spanned by the dominant eigenvector. Precisely, $A^{l}\to uv^{\top}$ as $l\rightarrow\infty$, where $Au=u$ and $A^{\top}v=v$ [40].
As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ converges in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$. Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{\top}X^{(l)}\|$, converges to zero.
@@ 3.1 Existing Oversmoothing Metrics
Motivated by the discussion about the linear case, oversmoothing is thus typically quantified and analysed in terms of the convergence of some node similarity metrics towards zero. In particular, in most cases, it is measured exactly by the alignment of the features with the dominant eigenvector of the matrix $A$. The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features [4, 33]
$E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2},$ (1)
where $u_{i}$ is the $i$-th entry of the dominant eigenvector of the message-passing matrix. It thus immediately follows from our discussion on the linear setting that $E_{\mathrm{Dir}}(X^{(l)})$ converges to zero as $l\to\infty$ for a linear GCN with converging weights product $W^{(0)}\cdots W^{(l)}$. This intuition suggests that a similar behaviour may occur for “smooth-enough” nonlinearities.
In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and [4] have proved that, using LeakyReLU activation functions, it holds $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
Similarly, in the case of GATs, the matrices $A_{i}$ are all row stochastic, meaning that $u_{i}=1$ for all $i$. In this case, it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero [45]
@@ 3.2 A Unifying Perspective Based on the Eigenvectors of Nonlinear Activations
We say that a vector $u\in\mathbb{R}^{N}\setminus\{0\}$ is an eigenvector of the (nonlinear) activation function $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{n}$ if for any $t\in\mathbb{R}\setminus\{0\}$, there exists $\mu_{t}\in\mathbb{R}$ such that $\sigma(tu)=\mu_{t}u$.
@@
This assumption recurs throughout our theoretical analysis, aligning with existing results in the literature. For example, (a) the proof by [4] applies to GCNs with LeakyReLU, where the dominant eigenvector of $A$ is nonnegative by the Perron-Frobenius theorem and, therefore, an eigenvector of LeakyReLU; and (b) the proof by [45] holds for stochastic message-passing matrices $A^{(l)}$, which inherently share a common dominant eigenvector with constant entries, that is an eigenvector for any pointwise nonlinear activation function.
@@
Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$, $l=1,\dots,L$, be a GNN such that $u$ is the dominant eigenvector of $A_{l}$ for any $l$ and is also an eigenvector of the activation function $\sigma$. If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, and $\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0$, then
$E_{\mathrm{Proj}}(X^{(L)})\to 0\quad\text{ as }\quad L\to\infty.$
Note, in particular, that in the case of GCNs, the matrix $A^{(l)}=A$ is symmetric, and thus $\|I-\mathcal{P}A^{(l)}\|_{2}=\lambda_{2}$. Therefore, when $\sigma=\text{LeakyReLU}$, we obtain the result by [4] as convergence to zero is guaranteed if $\|W^{(l)}\|_{2}\leq\lambda_{2}$. Recall that, as discussed above, the choice $\sigma=\text{LeakyReLU}$ satisfies our eigenvector assumption since $u\geq 0$ by the Perron-Frobenius theorem, and thus LeakyReLU$(tu)=\alpha tu$ with $\alpha$ depending only on the sign of $t$. Similarly, in the case of GATs, the matrices $A^{(l)}$ are stochastic for all $l$, implying that $u=\mathbbm{1}$ is the constant vector with $(u)_{i}=1$ for all $i$. If $\sigma=\otimes\psi$ is a nonlinear activation function acting entry-wise through $\psi$, then $\sigma(t\mathbbm{1})=\psi(t)\mathbbm{1}$. Therefore, Theorem 3.2 implies that if the weights are sufficiently small, the features align independently of the activation function used. This is consistent with the results in [45]. However, we note that the bounds on the weights required by Theorem 3.2 and those in [45] on the weights $W^{(l)}$ are not identical, and it is unclear which of the two is more significant. Nonetheless, in both cases, having bounded weights along with any $1$-Lipschitz pointwise activation function is a sufficient condition for observing oversmoothing via $E_{\mathrm{Proj}}$ in GATs. In addition to offering a different and unifying theoretical perspective on the results in [4, 45], we highlight the simplicity of our eigenvector-based proof, which we hope provides added clarity on this phenomenon.